# Bridging Binarization: Causal Inference with Dichotomized Continuous Exposures

### v4

@@ A.2 Inverse Probability Weighting (IPW) Estimator
The IPW estimator $\widehat{\Psi}^{IPW}$ requires an estimator of the propensity score $\pi_{\mathcal{A}}(w)=Pr(T=1|W=w)$ for all $w\in\text{supp}(W)$. If this estimate is consistent for the true propensity score, then the estimator will be consistent. Denote this consistent estimate $\widehat{\pi_{\mathcal{A}}}(w)$. Then, the IPW estimators for $\Psi_{BATE}$ and $\Psi_{CAB}$ are given by
@@ A.3 Augmented IPW Estimator (AIPW)
The AIPW estimator $\widehat{\Psi}^{AIPW}$ requires both an estimator of the conditional mean $m(t,w)\equiv E[Y|t,w]$ and an estimator of the propensity score conditional on confounders $\pi_{\mathcal{A}}(w)$. Denote these estimates $\widehat{m}(t,w)$ and $\widehat{\pi_{\mathcal{A}}}(w)$, respectively. Researchers can use a variety of machine learning (ML) algorithms to create these estimates, providing more flexibility. Super learning can also be used to determine the optimal predictor (or ensemble of predictors) from a class of many candidate estimators based on the data [34]. Researchers should employ sample splitting and cross fitting to reduce over-fitting and ensure that the estimates converge to the truth sufficiently quickly [35, 36].
@@
Then, the variance estimates for $\widehat{\Psi}_{BATE}^{AIPW}$ and $\widehat{\Psi}_{CAB}^{AIPW}$ can be calculated as
$\widehat{\mathbb{V}}\left(\widehat{\Psi}_{BATE}^{AIPW}\right)$ $=\frac{1}{n}{\sum}_{i=1}^{n}\widehat{\phi}(O_{i})_{BATE}^{2}$
$\widehat{\mathbb{V}}\left(\widehat{\Psi}_{CAB}^{AIPW}\right)$ $=\frac{1}{n}{\sum}_{i=1}^{n}\widehat{\phi}(O_{i})_{CAB}^{2}.$
A.4 Targeted Maxmimum Likelihood Estimator (TMLE)
The TMLE estimator $\widehat{\Psi}^{TMLE}$ is a semi-parametrically efficient, doubly-robust estimator. We provide a brief overview of the steps required below. For a more comprehensive introduction to TMLE for applied work in the binary exposure setting, we refer researchers to the 2017 paper “Targeted Maximum Likelihood Estimation for Causal Inference in Observational Studies” by Schuler and Rose [37]. The TMLE presented for the BATE is equivalent to the binTMLE presented in [20] and achieves the semi-parametric efficiency bound under the assumption of missingness at random.
@@ A.4 Targeted Maxmimum Likelihood Estimator (TMLE)
    • If estimating $\Psi_{BATE}$, $H=\frac{1(T=1)}{g_{0}(W)}-\frac{1(T=0)}{1-g_{0}(W)}$
    • If estimating $\Psi_{CAB}$, $H=1-\frac{1(T=1)}{g_{0}(W)}$ when considering $E[Y_{\tilde{A}_{1}}]$ and $H=1-\frac{1(T=0)}{1-g_{0}(W)}$ when considering $E[Y_{\tilde{A}_{0}}]$.
@@
    $\text{logit}(Y)=\text{logit}(\hat{Y})+\beta H(T,W)$
@@
    $\text{logit}(\hat{Y}_{\tilde{A}_{0}}^{*})$ $=\text{logit}(\widehat{Y}_{\tilde{A}_{0}})+\hat{\beta}\frac{1}{1-g_{0}(W)}$
@@
    $\hat{\Psi}_{BATE}^{TMLE}$ $=\frac{1}{n}\sum_{i=1}^{n}\hat{Y}_{\tilde{A}_{1}}^{*}-\hat{Y}_{\tilde{A}_{0}}^{*}$ (17)
    $\hat{\Psi}_{CAB}^{TMLE}$ $=\frac{1}{n}\sum_{i=1}^{n}\hat{Y}^{*}-\hat{Y}_{\tilde{A}_{t}}^{*}$ (18)
To estimate the variance, researchers can either calculate influence curve based estimates or use the nonparametric bootstrap. The influence curve based estimates are the same as the AIPW estimator as in Appendix A.2 and have the same consistently conditions.
@@ Appendix B When is the regression estimator consistent?
Consider a cut-off at $A=a_{0}$ such that $T=1$ when $A\geq a_{0}$ and $T=0$ when $A<a_{0}$. Then, we can consider the expectation of $Y$ given $T=1$ and $W$.
@@
$=\frac{1}{\pi_{\mathcal{A}}(w)}\left(\beta_{0}+\beta_{W}W\right)+\beta_{A}E[g_{A}(\tilde{A}_{1})|W]+\beta_{int}WE[g_{int}(\tilde{A}_{1})|W]$
Similarly for $E[Y|T=0,W]$, we have
$E[Y|T=0,W]=\frac{1}{1-\pi_{\mathcal{A}}(w)}\left(\beta_{0}+\beta_{W}W\right)+\beta_{A}E[g_{A}(\tilde{A}_{0})|W]+\beta_{int}WE[g_{int}(\tilde{A}_{0})|W]$

### v5

@@ A.2 Inverse Probability Weighting (IPW) Estimator
The IPW estimator $\widehat{\Psi}^{IPW}$ requires an estimator of the propensity score $\pi_{\mathcal{A}}(w)=Pr(T=1|W=w)$ for all $w\in\text{supp}(W)$. If this estimator is consistent for the true propensity score, then the IPW estimator will be consistent. Denote this consistent estimate $\widehat{\pi_{\mathcal{A}}}(w)$. Then, the IPW estimators for $\Psi_{BATE}$ and $\Psi_{CAB}$ are given by
@@ A.3 Augmented IPW Estimator (AIPW)
The AIPW estimator $\widehat{\Psi}^{AIPW}$ is a semi-parametrically efficient, doubly-robust estimator. $\widehat{\Psi}^{AIPW}$ requires both an estimator of the conditional mean $m(t,w)\equiv E[Y|t,w]$ and an estimator of the propensity score conditional on confounders $\pi_{\mathcal{A}}(w)$. Denote these estimates $\widehat{m}(t,w)$ and $\widehat{\pi_{\mathcal{A}}}(w)$, respectively. Researchers can use a variety of machine learning (ML) algorithms to create these estimates, providing more flexibility. Super learning can also be used to determine the optimal predictor (or ensemble of predictors) from a class of many candidate estimators based on the data [20]. Researchers should employ sample splitting and cross fitting to reduce over-fitting and ensure that the estimates converge to the truth sufficiently quickly [33, 4].
@@
Then, the variance estimates for $\widehat{\Psi}_{BATE}^{AIPW}$ and $\widehat{\Psi}_{CAB}^{AIPW}$ can be calculated as the estimated variance of the influence curve divided by the sample size $n$:
$\widehat{\mathbb{V}}\left(\widehat{\Psi}_{BATE}^{AIPW}\right)$ $=\frac{1}{n}\widehat{\text{Var}}\left(\widehat{\phi}(O_{i})_{BATE}\right)$
$\widehat{\mathbb{V}}\left(\widehat{\Psi}_{CAB}^{AIPW}\right)$ $=\frac{1}{n}\widehat{\text{Var}}\left(\widehat{\phi}(O_{i})_{CAB}\right).$
A.4 Targeted Maximum Likelihood Estimator (TMLE)
The TMLE estimator $\widehat{\Psi}^{TMLE}$ is another semi-parametrically efficient, doubly-robust estimator. For a more comprehensive introduction to TMLE for applied work in the binary exposure setting, we refer researchers to the 2017 paper “Targeted Maximum Likelihood Estimation for Causal Inference in Observational Studies” by Schuler and Rose [27]. The TMLE presented for the BATE is equivalent to the binTMLE presented in [32] and achieves the semi-parametric efficiency bound under the assumption of missingness at random. We provide a brief overview of the steps required below.
@@ A.4 Targeted Maximum Likelihood Estimator (TMLE)
    • If estimating $\Psi_{BATE}$, $H=\frac{1(T=1)}{g_{0}(W)}-\frac{1(T=0)}{1-g_{0}(W)}$.
    • If estimating $\Psi_{CAB}$, $H=\frac{1(T=1)}{g_{0}(W)}-1$ when considering $E[Y_{\tilde{A}_{1}}]$ and $H=\frac{1(T=0)}{1-g_{0}(W)}-1$ when considering $E[Y_{\tilde{A}_{0}}]$.
@@
    $\text{logit}(Y)=\text{logit}(\hat{Y})+\beta H(T,W).$
@@
    $\text{logit}(\hat{Y}_{\tilde{A}_{0}}^{*})$ $=\text{logit}(\widehat{Y}_{\tilde{A}_{0}})+\hat{\beta}\frac{1}{1-g_{0}(W)}.$
@@
    $\hat{\Psi}_{BATE}^{TMLE}$ $=\frac{1}{n}\sum_{i=1}^{n}\hat{Y}_{\tilde{A}_{1}}^{*}-\hat{Y}_{\tilde{A}_{0}}^{*}$
    $\hat{\Psi}_{CAB}^{TMLE}$ $=\frac{1}{n}\sum_{i=1}^{n}\hat{Y}_{\tilde{A}_{t}}^{*}-\hat{Y}^{*}.$
To estimate the variance, researchers can either calculate influence curve based estimates or use the nonparametric bootstrap. The influence curve based estimates are the same as the AIPW estimator as in Appendix A.2 and have the same consistency conditions.
@@ Appendix B When is the regression estimator consistent?
Consider a cut-off at $A=a_{0}$ such that $T=1$ when $A\geq a_{0}$ and $T=0$ when $A<a_{0}$. Then, we can consider the expectation of $Y$ given $T=1$ and $W$:
@@
$=\frac{1}{\pi_{\mathcal{A}}(w)}\left(\beta_{0}+\beta_{W}W\right)+\beta_{A}E[g_{A}(\tilde{A}_{1})|W]+\beta_{int}WE[g_{int}(\tilde{A}_{1})|W].$
Similarly, for $E[Y|T=0,W]$, we have
$E[Y|T=0,W]=\frac{1}{1-\pi_{\mathcal{A}}(w)}\left(\beta_{0}+\beta_{W}W\right)+\beta_{A}E[g_{A}(\tilde{A}_{0})|W]+\beta_{int}WE[g_{int}(\tilde{A}_{0})|W].$